### v2

# Persistent coherent quantum dynamics in 2D long-range magnets via magnon binding

Abstract
The dynamics of 2D long-range quantum magnets represents a current frontier in experimental physics such as in Rydberg atomic systems or trapped ions. In this work we address theoretical challenges in understanding these dynamics by combining large-scale neural quantum state simulations with an effective theory. Our findings uncover a mechanism for persistent coherent quantum dynamics and slow relaxation in 2D long-range quantum magnets. Demonstrated on the 2D transverse-field quantum Ising model with power-law decaying interactions, we observe long-lived oscillatory behavior after quenching the system from a ferromagnetic product state. We explain this phenomenon by the formation of magnon bound states, generated by effective attractive long-range magnon interactions. Our results highlight a generic mechanism for long-lived quantum coherence in 2D quantum magnets that can be directly observed in current quantum simulation platforms.
Introduction.
Figure 1: Slow dynamics and magnon binding. (A) Spatiotemporal normalized connected correlator $\tilde{C}(d,t)$ after a quench from $|\!\downarrow\cdots\downarrow\rangle$ on an $11\times 11$ lattice ($\alpha=3$, $g/J=0.2$), showing slow dynamics and oscillations. (B) Schematic illustration of two magnons at separation $d$. The lower panel sketches the effective attractive potential $U(d)$ for $\alpha=3$ from Eq. 4. Inset: the corresponding two-magnon bound-state probability peaked at $\tilde{r}=\sqrt{5}$ on an $101\times 101$ lattice.
Figure 2: Fourier spectroscopy for the post-quench dynamics ($9\times 9$, $\alpha=3$ and $g/J=0.5$). (A) Spatiotemporal connected correlator $\tilde{C}(d,t)$, together with the site-averaged longitudinal magnetization $\langle\hat{S}_{i}^{z}(t)\rangle$. (B) FFTs $\mathcal{F}[\,C(d,t)\,]$ for $d=1,2,3$, and $\mathcal{F}[\,\langle\hat{S}_{i}^{z}(t)\rangle\,]$, plotted versus frequency $\omega/J$. Vertical dashed lines mark energy gaps $\Delta^{(\nu\leftrightarrow\nu^{\prime})}_{i,j}=E^{\nu}_{i}-E^{\nu^{\prime}}_{j}$ from effective Hamiltonians restricted to the $\nu$ magnon sectors. The matching between spectral peaks and energy gaps quantitatively validates the few-body effective description. Inset: Fourier spectra of exact dynamics for a $5\times 5$ system (evolved up to $Jt=200$).
Long-range quantum magnets arise across atomic, molecular, optical, and solid-state platforms [15], with prototypical realizations provided by spin models with power-law couplings in trapped ions [7, 30, 21, 5, 41, 27, 23], Rydberg atom arrays [8, 16, 6], and dipolar quantum gases [38, 19, 13, 14, 1]. In 1D, theoretical and experimental studies have established that long-range interactions modify correlation spreading and the effective light-cone structure [18, 30], affect entanglement growth [21], and generate prethermal plateaus and long-lived oscillations that delay thermalization [25]. These phenomena have been linked to emergent quasiparticles, such as domain walls or magnons, their mutual interactions, and the bound states they form [4, 37, 21, 23, 24, 17, 26].
However, extending these insights to two dimensions remains a central open problem. The exponential growth of the Hilbert space together with the rapid buildup of entanglement makes well-established methods such as exact diagonalization and standard tensor-network methods (e.g., tDMRG/MPS) challenging. Consequently, the quantum dynamics of 2D long-range magnets remain largely unexplored. Beyond the numerical challenge, the interplay of 2D geometry and power-law interactions can give rise to dynamical mechanisms that are distinct from both the one-dimensional case and the nearest-neighbor limit, raising basic questions: What are the relevant elementary excitations? Can nonlocal interactions in 2D bind excitations over several lattice sites? How does this impact thermalization?
In this Letter, we address these questions by combining large-scale neural quantum state (NQS) simulations [9, 31, 33, 32] with an effective theoretical description. We uncover a generic mechanism for persistent coherent quantum dynamics and slow relaxation in 2D long-range quantum magnets. As an experimentally relevant benchmark, we focus on the 2D transverse-field Ising model with power-law interactions, accessing system sizes and evolution times beyond established numerical approaches. Quenching from a fully polarized ferromagnetic product state, we observe slow relaxation characterized by persistent, long-lived oscillations, a striking departure from the general expectation of rapid thermalization. To interpret this behavior, we construct an effective theory whose eigenspectrum shows excellent agreement with the Fourier spectral analysis of the real-time dynamics. Crucially, the effective model reveals that long-range interactions mediate a power-law attraction between magnons, yielding multi-magnon bound states even at separations of several lattice sites and producing a large manifold of bound states. We thus identify extended magnon bound states as the quasiparticles governing slow relaxation in the 2D long-range Ising model, and argue that magnons form bound states whenever long-range attractive interactions compete with weak kinetic terms, making magnon binding a generic mechanism for persistent coherent dynamics in 2D long-range magnets.
Model.— We consider the 2D long-range quantum Ising model described by the Hamiltonian
$\hat{H}=-\frac{J}{\mathcal{N}_{\alpha}}\sum_{i\neq j}\frac{\hat{S}_{i}^{z}\hat{S}_{j}^{z}}{r_{ij}^{\alpha}}-g\sum_{i}\hat{S}_{i}^{x},$ (1)
where $\hat{S}_{i}^{\mu}$ ($\mu=x,y,z$) are spin-$1/2$ operators on a square lattice of size $L\times L$ with periodic boundary conditions, and $r_{ij}$ denotes the distance between sites $i$ and $j$. The exponent $\alpha$ controls the range of the interaction: small values of $\alpha$ correspond to slowly decaying couplings, while $\alpha\to\infty$ recovers the nearest-neighbor limit. We employ Kac normalization $\mathcal{N}_{\alpha}=(L^{2}-1)^{-1}\sum_{{i<j}}r_{ij}^{-\alpha}$, which keeps the energy per site finite as $L\rightarrow\infty$.
Figure 3: Eigenstate structure in the two-magnon sector ($101\times 101$, $g/J=0.2$). Inverse participation ratio (IPR) versus mean pair separation $\bar{d}$ for all eigenstates. (A) $\alpha=2$: a continuous tail to large $\bar{d}$ indicates bound states and quasilocalized states. (B) $\alpha=3$: intermediate regime with bound states at intermediate separations. (C) $\alpha=6$: sharp drop in IPR with $\bar{d}$ signals short-range binding only. Insets: real-space distributions of representative eigenstates in the $(\mathbf{r}_{1}-\mathbf{r}_{2})$ plane.
[…]
$C(d,t)=\langle\hat{S}^{z}_{i}(t)\hat{S}^{z}_{i+d}(t)\rangle-\langle\hat{S}^{z}_{i}(t)\rangle\langle\hat{S}^{z}_{i+d}(t)\rangle,$ (2)
measured along the x-direction of the lattice. For comparison across different $\alpha$ and $g$, we plot the rescaled correlator $\tilde{C}(d,t)=C(d,t)\cdot\frac{8\mathcal{N}_{\alpha}J^{2}}{g^{2}}$, which provides a suitable normalization across multiple parameter regimes.
To reach system sizes beyond the limits of exact diagonalization for real-time simulations, we employ neural quantum states (NQS), a variational representation of quantum wavefunctions using artificial neural networks. Within NQS, the time-dependent state $|\psi_{\bm{\theta(t)}}\rangle=\sum_{S}\psi_{\bm{\theta(t)}}(S)|S\rangle$, with $|S\rangle$ denoting the full set of $2^{N}$ spin configurations, is represented through wave-function amplitudes $\psi_{\bm{\theta}(t)}(S)$ which are parametrized by network weights $\bm{\theta}$. The evolution of these parameters, $\bm{\theta}(t)$, is governed by the time-dependent variational principle. Importantly, NQS provides a numerically exact solution of the dynamics in the sense that convergence is guaranteed upon increasing the size of the underlying artificial neural network. Specifically, we use a convolutional neural network (CNN) architecture based on the ResNet framework Ref. [20, 11, 12], which is optimized for translational invariance and to capture 2D spatial correlations.
Our numerical results, obtained via NQS simulations, show that for an $11\times 11$ lattice [Fig. 1(A)], the spatiotemporal correlations $\tilde{C}(d,t)$ build up only slowly, exhibiting persistent oscillations at the same time. Extending the dynamics to times up to $Jt=60$ for a $9\times 9$ lattice [Fig. 2(A)] reveals that the oscillations in both the longitudinal magnetization $\langle\hat{S}_{i}^{z}(t)\rangle$ and the correlation $\tilde{C}(d,t)$ are unconventionally long-lived. To resolve their characteristic frequencies, we perform a discrete Fourier transform $\mathcal{F}$ of $C(d,t)$ and $\langle\hat{S}_{i}^{z}(t)\rangle$, applying a Hamming window over the interval Jt=5Jt=56060. The resulting spectra exhibit sharp peaks [Fig. 2(B)], signaling the emergence of well-defined modes. To identify their microscopic origin and account for the slow dynamics, we next develop an effective low-energy theory.
[…]
$E_{\nu}(\{i_{m}\})-E_{0}=2\nu J\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\mathcal{N}_{\alpha}}\sum_{1\leq k<l\leq\nu}\frac{1}{r^{\alpha}_{i_{k}i_{l}}}.$
The first term, which grows linearly with the number of magnons, represents the cost of an isolated magnon. The second term reflects an attractive interaction energy between all magnon pairs.
In the nearest-neighbor limit ($\alpha=\infty$), the interaction term vanishes unless $r_{i_{k}i_{l}}=1$. The energy expression then reduces to a constant proportional to the number of domain-wall links, implying that the excitation spectrum is classified solely by the interface length of spin-flip domains. For finite $\alpha$, however, the long-range couplings generate a much richer structure. Specifically, the single-magnon sector forms the lowest excited band, followed by the two-magnon manifold. Higher sectors contain states with three or more magnons, but their ordering is highly sensitive to the geometry of how the magnons are arranged. For example, a compact four-magnon cluster lies energetically below a configuration of three well-separated magnons. Within this hierarchy, the states directly above the two-magnon manifold are those three-magnon configurations in which at least two magnons occupy nearest-neighbor sites.
Having established the structure of low-energy excitations in the pure Ising Hamiltonian, we now focus on the influence of the perturbation $V=-g\sum_{i}\hat{S}^{x}_{i}$. Since, for finite $\alpha$, the $\nu=0,1,2$ sectors together with the three-magnon sector containing at least a nearest-neighbor pair constitute well-separated excitation bands with gaps of $\mathcal{O}(J)$, the effect of a weak transverse field strength $g$ can be treated perturbatively. Applying a Schrieffer–Wolff (SWT) transformation up to second order in $g$ in each sector, we integrate out couplings to other sectors at leading order and thereby obtain an effective Hamiltonian acting solely within the sector with corrections to this structure arising only at order $\mathcal{O}(g^{3}/J^{2})$. For the two-magnon manifold, imposing translational symmetry and restricting to the zero-momentum subspace, yields the effective Hamiltonian
[…]
where the basis states $|\mathbf{d}\rangle$ represent two magnons separated by the relative coordinate $\mathbf{d}$. $E_{2}\approx E_{0}+4J-g^{2}L^{2}/8J$ fixes the overall energy of the two-magnon manifold. A central finding of our effective description is that two magnons experience an attractive interaction potential of long-range character,
$\ U(\mathbf{d})$ $=-\left(2J-\frac{g^{2}}{2J}\right)\frac{1}{\mathcal{N}_{\alpha}|\mathbf{d}|^{\alpha}}+\delta U(\mathbf{d}),$ (4)
with $\delta U(\mathbf{d})$ denoting corrections proportional to $g^{2}/J$, which at large distances $|\mathbf{d}|$ provide only a subleading contribution to the attractive interactions proportional to $\frac{-g^{2}}{8J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{2(\alpha-1)}}$, as long as $\alpha>2$. The off-diagonal amplitudes $t_{\mathbf{d},\mathbf{d}^{\prime}}$ are real and symmetric, describing correlated long-range pair hoppings, which at large $|\mathbf{d}-\mathbf{d^{\prime}}|$ are given by $t_{\mathbf{d},\mathbf{d}^{\prime}}\approx\frac{-g^{2}}{8J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}-\mathbf{d}^{\prime}|^{\alpha}}$ (explicit derivations are provided in the supplementary material). For all numerical calculations, however, we employ the full (non-asymptotic) effective Hamiltonian derived via the Schrieffer–Wolff transformation. Notice that while an attractive interaction potential might be naturally associated with the formation of bound states in the case of short-range kinetics [39, 10], we find that the interplay between the long-range attractive potential and the correlated pair long-range hoppings produces a rich spectrum of two-magnon bound states.
Figure 4: $\tilde{C}(d,t)$ for a $9\times 9$ lattice with $g/J=0.2$. The panels compare interaction exponents $\alpha=2,3$ with $\alpha=6$ and the nearest-neighbor limit $\alpha=\infty$.
To validate this theoretical framework, we diagonalize $\hat{H}_{\mathrm{eff}}^{(\nu)}$, and then compare the resulting level spacings with peak positions extracted from the Fourier spectra of $\langle\hat{S}^{z}_{i}(t)\rangle$ and $C(d,t)$. Figure 2(B) shows these spectra for $9\times 9$ lattice with $g/J=0.5$ and $\alpha=3$ overlaid with excitation gaps $\Delta^{(\nu\leftrightarrow\nu^{\prime})}_{i,j}=|E^{\nu}_{i}-E_{j}^{\nu^{\prime}}|$, where $E^{\nu}_{i}$ denotes the ithi-th energy level of $\hat{H}_{\mathrm{eff}}^{(\nu)}$. The dominant peak at $\omega/J\simeq 1.89$ corresponds to $\Delta^{(0\leftrightarrow 1)}_{1,1}$, i.e., creating a single magnon above the polarized background (see Supplementary Material for the single-magnon dispersion). Additional peaks arise from $\Delta^{(0\leftrightarrow 2)}$ and $\Delta^{(1\leftrightarrow 2)}$, corresponding to creating a two-magnon bound state out of the polarized state or out of the single-magnon sector, respectively. We further resolve a peak associated with a three-magnon process $\Delta^{(1\leftrightarrow 3)}$, corresponding to the creation of three tightly bound magnons. Some peaks appear slightly offset from the exact gap values due to the limited frequency resolution imposed by the NQS time window. As a result, the discrete frequency grid does not include points exactly at those gaps, so the spectral weight appears at the nearest available frequency bin, leading to a small apparent mismatch.
To achieve a more refined frequency resolution, the inset displays the FFT of the exact dynamics for a $5\times 5$ system evolved up to $Jt=200$. The resulting spectrum agrees very well with the effective predictions, except for the red line. This deviation arises because, for $L\leq 5$, the excitations above the two-magnon manifold, those involving three or more magnons, do not form a well-separated band, which limits the applicability of the SWT description. Nevertheless, exact diagonalization reveals that this peak originates from a compact four-magnon cluster. These observations show that the slow dynamics are controlled by a sparse set of well-defined few-magnon excitations rather than a broad many-body continuum. Furthermore, the clear one-to-one correspondence between the spectral peaks extracted from the NQS dynamics and the energy gaps of the effective theory confirms the accuracy of both approaches up to at least $g/J=0.5$.
We emphasize that the effective Hamiltonian is broadly applicable to 2D long-range magnets. Whenever a microscopic model features power-law decaying ferromagnetic couplings, the low-energy sector above a fully polarized ferromagnetic state is governed by few-magnon excitations. Adding kinetic terms such as spin flips, hopping, or pair creation and annihilation, or any combination thereof, yields, under a SW transformation, an effective Hamiltonian of very similar structure.
Emergence of magnon bound states.—Having validated the effective theory with our exact numerical simulations with NQS, we now aim to analyze in more detail the two-magnon eigenstates. In this section, we address two central questions: (i) whether the long-range attractive potential binds two magnons beyond the nearest-neighbor separation, (ii) how the spatial structure of these states depends on the interaction exponent $\alpha$. To characterize these states, we analyze the eigenstates $|\psi(\mathbf{d})\rangle$ of the effective two-magnon Hamiltonian $\hat{H}_{\text{eff}}^{(2)}$. For each eigenstate, we compute its inverse participation ratio (IPR), $\mathrm{IPR}[\psi]=\sum_{\mathbf{d}}|\psi(\mathbf{d})|^{4}$, which distinguishes localized bound pairs (high IPR) from delocalized scattering states (low IPR). Intermediate IPR values correspond to quasilocalized states, whose wavefunctions contain a localized core accompanied by a weak delocalized tail. We also compute the mean pair separation, $\bar{d}=\sum_{\mathbf{d}}|\psi(\mathbf{d})|^{2}\,|\mathbf{d}|$, to determine the average relative positions. Figure 3 shows the resulting IPR-$\bar{d}$ distribution for several power-law exponents $\alpha$ on a $101\times 101$ lattice with $g/J=0.2$, together with representative real-space probability densities in the relative coordinate $(\mathbf{r}_{1}-\mathbf{r}_{2})$.
For strongly long-ranged interactions ($\alpha=2$), many eigenstates exhibit a large IPR for mean separations up to $\bar{d}=4$, followed by a continuous decay of IPR with increasing $\bar{d}$ [Fig. 3(A)]. This indicates that the spectrum contains bound magnon pairs with separations as large as $d=4$. Furthermore, a substantial fraction of eigenstates displays quasilocalized character, with intermediate values of both IPR and $\bar{d}$, with a localized component that persists out to separations of $d=6$. For $\alpha=3$, the number of quasilocalized states is noticeably reduced [Fig. 3(B)], although bound pairs still survive up to $\bar{d}\approx 4$. These findings demonstrate that long-range interactions stabilize magnons at distances well beyond nearest neighbors and also enhance the number of localized and quasilocalized states. By contrast, in the short-range regime ($\alpha=6$), only a small number of eigenstates exhibit high IPR, while the majority show low IPR at large $\bar{d}$. Consequently, the magnon binding is restricted to short separations up to $d\approx\sqrt{2}$ [Fig. 3(C)].
Notice that within the approximation, the effective interaction $U(\mathbf{d})$ in Eq. 4 becomes less attractive as $g$ increases, which correspondingly suppresses the formation of bound states at larger separations. For example, at $g/J=0.5$ we find that for $\alpha=2$ and $\alpha=3$, localized states persist up to separations $d=2$ and $d=\sqrt{2}$, respectively. For $\alpha=6$ and $\alpha=\infty$, only nearest-neighbor bound states remain.
Figure 4 shows the time evolution of the correlations $\tilde{C}(d,t)$ for various interaction exponents $\alpha$ at $g/J=0.2$. A clear distinction emerges between the short-range regime ($\alpha=6,\infty$) and the long-range regime ($\alpha=2,3$). For large $\alpha$, the signal is stronger at the nearest-neighbor distance $d=1$, and beyond this separation, the dynamics becomes nearly monochromatic, consistent with the presence of only tight bound magnon pairs and a few characteristic oscillation frequencies. In contrast, for small $\alpha$, the correlator exhibits multiple frequency components and a significant signal at $d>1$, reflecting the existence of several extended bound states.
Discussion and Outlook. Our NQS simulations and effective theory establish magnon binding as a generic microscopic mechanism for persistent coherent dynamics and slow relaxation in 2D long-range magnets. Future experiments with Rydberg arrays and trapped ions could directly probe these spatially extended bound magnons using site-resolved correlation spectroscopy, enabling real-time visualization of bound-pair formation and decay. This mechanism is fundamentally distinct from the 1D case, where domain walls are the elementary excitations and long-range interactions confine domain-wall kinks [25], and from the 2D nearest-neighbor limit, where slow dynamics has been linked to approximate Hilbert-space fragmentation due to the near-conservation of domain-wall length in the weak-field regime [40, 2, 3, 35, 29]. Beyond the transverse-field Ising model studied here, our framework, in principle, applies broadly to 2D long-range magnets with diverse kinetic processes. Accordingly, we expect that magnon binding could be a generic feature of 2D long-range quantum magnets, potentially also in 3D, where magnons again constitute the elementary excitations.
Data availability.—Data to reproduce all figures are available at Zenodo [28].
Acknowledgements.—The NQS dynamics is simulated using the jVMC package [34], and the exact dynamics are performed with the QuSpin package [36]. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 853443). This work was supported by the German Research Foundation (DFG) via project 492547816 (TRR 360). We gratefully acknowledge the scientific support and high-performance computing resources provided by the LiCCA HPC cluster at the University of Augsburg and by the Erlangen National High Performance Computing Center (NHR) of the Friedrich-Alexander-Universität Erlangen–Nürnberg (NHR Project No. nqsQuMat). The LiCCA cluster is co-funded by the DFG (Project ID 499211671). NHR funding is provided by federal and Bavarian state authorities, and the NHR@FAU hardware is partially funded by the DFG (Grant No. 440719683).
Note added.—During the preparation of this manuscript, we became aware of related work by Kaltenmark et al. [22], who study spectroscopic signatures in a kinetically constrained long-range interacting two-dimensional spin system. While their microscopic model and parameter regime differ from ours, both works highlight the emergence of nontrivial elementary excitations in long-range interacting quantum systems.
Supplemental Material: Persistent coherent quantum dynamics in 2D long-range magnets via magnon binding
Appendix A Schrieffer-Wolff transformation.
This section derives the effective few-magnon Hamiltonian using the Schrieffer–Wolff (SW) transformation. A configuration with $\nu$ magnons at positions $i_{1},\ldots,i_{\nu}$} has unperturbed energy
$E_{\nu}(\{i_{m}\})=E_{0}+2\nu J\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\mathcal{N}_{\alpha}}\sum_{1\leq k<l\leq\nu}\frac{1}{r^{\alpha}_{i_{k}i_{l}}}.$
where $E_{0}=-(L^{2}-1)/2$ denotes the energy of the fully polarized state and only energy differences enter the SW expansion. With zero magnon states being the lowest energy band, $E_{\nu}(\{i_{m}\})$ helps to determine the structure of low-energy excitations. In the nearest-neighbor limit ($\alpha=\infty$), the energy levels are classified solely by the interface length. For finite $\alpha$, zero-, one-, and two-magnon sectors form well-separated excitation bands with gaps of order $\mathcal{O}(J)$.
Introducing a weak transverse field $g\hat{T}=-\sum_{i}\hat{S}_{i}^{x}$, couples different $\nu$ sectors only perturbatively. To eliminate these couplings at leading order, we perform the Schrieffer–Wolff (SW) transformation $\hat{H}_{\rm SW}=e^{\hat{S}}\hat{H}e^{-\hat{S}}$, and choose the generator $\hat{S}$ such that $\hat{S},\hat{H}_{0}]=-g\,\hat{T}$, which removes matrix elements between the $\nu$ and $\nu\!\pm\!1$ sectors to first order in $g$. Projecting onto the $\nu$-magnon manifold $\mathcal{P}_{\nu}$ gives
$\hat{H}_{\rm eff}^{(\nu)}=P_{\nu}\hat{H}_{0}P_{\nu}+\frac{g^{2}}{2}\,P_{\nu}[\hat{S},\hat{T}]P_{\nu}+\mathcal{O}(g^{3}/J^{2}).$
Using the standard expression for $\hat{S}$ in terms of virtual transitions, we obtain
$\hat{H}_{\rm eff}^{(\nu)}=\sum_{m}E_{\nu}(m)\,|m\rangle\!\langle m|+\frac{g^{2}}{2}\sum_{m,n,\beta}\langle m|\hat{T}|\beta\rangle\,\langle\beta|\hat{T}|n\rangle\left(\frac{1}{E_{m}-E_{\beta}}+\frac{1}{E_{n}-E_{\beta}}\right)|m\rangle\!\langle n|,$
where $\{|m\rangle\}$ spans the $\nu$-magnon space $\mathcal{P}_{\nu}$, and $|\beta\rangle$ runs over virtual states in the complement $\mathcal{Q}_{\nu}=1-\mathcal{P}_{\nu}$ (i.e., the $\nu-1$ and $\nu+1$ sectors). This equation provides the leading $\mathcal{O}(g^{2}/J)$ corrections that define the effective few-magnon Hamiltonian used in the main text.
Appendix B Effective zero–magnon Hamiltonian
In the zero–magnon sector ($\nu=0$) there is only a single basis state. The effective Hamiltonian $\hat{H}_{\rm eff}^{(0)}$ therefore reduces to the diagonal energy
$E_{\rm eff}^{(0)}=E_{0}+g^{2}\sum_{\beta}\frac{|\langle 0|\hat{T}|\beta\rangle|^{2}}{E_{0}-E_{\beta}},$
where $|\beta\rangle$ runs over all single–magnon states, and all single–magnon states lie at energy $E_{\beta}=E_{0}+2J\!\left(1-\frac{1}{L^{2}}\right).$ Summing over the $L^{2}$ possible single–flip configurations gives
$\boxed{E_{\rm eff}^{(0)}=E_{0}-\frac{g^{2}}{8J}\frac{L^{2}}{(1-1/L^{2})}\approx E_{0}-\frac{g^{2}L^{2}}{8J}.}$
Appendix C Effective single-magnon Hamiltonian.
In the single–magnon sector ($\nu=1$), the effective Hamiltonian $\hat{H}_{\rm eff}^{(1)}$ can be written as
[…]
Figure S1: Single-magnon dispersion for different power-law exponents $\alpha$. Left: dispersion $E_{1}(k_{x},k_{y}{=}0)-E_{0}$ along the $k_{x}$ direction. Right: dispersion along the high-symmetry path $X\to M\to\Gamma\to X\to S$. Long-range interactions induce a nonanalytic cusp at $\Gamma$, which gradually disappears as $\alpha$ becomes large.
Diagonal term.
The diagonal term also acquires a second–order correction in $g$ to $E_{1}$,
[…]
$=E_{0}+2J\left(1-\frac{1}{L^{2}}\right)+\frac{g^{2}}{8J\left(1-\frac{1}{L^{2}}\right)}-\frac{g^{2}}{4}\sum_{j\neq i}\frac{1}{2J\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\mathcal{N}_{\alpha}\,|r_{ij}|^{\alpha}}}.$
Taking the large–$L$ limit, and writing $\mathbf{d}=r_{ij}$, this simplifies to
$U\approx E_{0}+2J+\frac{g^{2}}{8J}-\frac{g^{2}}{8J}\sum_{|\mathbf{d}|\neq 0}\left(1-\frac{1}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}\right)^{-1}.$
Expanding $\left(1-\frac{1}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}\right)^{-1}$ to first order in $\frac{1}{|\mathbf{d}|^{\alpha}}$, we obtain
$U$ $\approx E_{0}+2J+\frac{g^{2}}{8J}-\frac{g^{2}}{8J}\sum_{|\mathbf{d}|\neq 0}\left(1+\frac{1}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}\right)$
$=E_{0}+2J+\frac{g^{2}}{8J}-\frac{g^{2}(L^{2}-1)}{8J}-\frac{g^{2}}{8J\mathcal{N}_{\alpha}}\sum_{|\mathbf{d}|\neq 0}\frac{1}{|\mathbf{d}|^{\alpha}}$
$=E_{0}+2J-\frac{g^{2}L^{2}}{8J}.$
Therefore, in the large–$L$ limit the diagonal part of $\hat{H}_{\rm eff}^{(1)}$ is well approximated by
$U\approx E_{0}+2J-\frac{g^{2}L^{2}}{8J}.$
Off–diagonal term.
The off–diagonal matrix elements between two single–magnon states are given by
[…]
$=\frac{g^{2}}{4}\left(\frac{1}{2J\left(1-\frac{1}{L^{2}}\right)}-\frac{1}{2J\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\,|r_{ij}|^{\alpha}}}\right).$
Taking again the large–$L$ limit and expanding to first order in $\frac{1}{|r_{ij}|^{\alpha}}$, we obtain
$t_{i,j}\approx-\frac{g^{2}}{8J\mathcal{N}_{\alpha}}\left(\frac{1}{|r_{ij}|^{\alpha}-1/\mathcal{N}_{\alpha}}\right)\approx-\frac{g^{2}}{8J\mathcal{N}_{\alpha}}\,\frac{1}{|r_{ij}|^{\alpha}}.$
Combining the diagonal and off–diagonal processes, the effective single–magnon Hamiltonian becomes
$|.\displaystyle\boxed{\hat{H}_{\rm eff}^{(1)}=\left(E_{0}+2J-\frac{g^{2}L^{2}}{8J}\right)\;-\;\frac{g^{2}}{8J\mathcal{N}_{\alpha}}\sum_{i\neq j}\frac{1}{|r_{ij}|^{\alpha}}\ket{i}\!\bra{j}.}$
Figure S1 displays the single-magnon dispersion for different power-law exponents.
Appendix D Effective two-magnon Hamiltonian.
In the two–magnon sector ($\nu=2$), the effective Hamiltonian $\hat{H}_{\rm eff}^{(2)}$ can be written as
[…]
$|.\displaystyle\quad\quad\quad\quad\quad\quad\quad\times\Bigl(\frac{1}{E_{2}(n_{1},n_{2})-E_{\beta}}+\frac{1}{E_{2}(m_{1},m_{2})-E_{\beta}}\Bigr)\ket{n_{1},n_{2}}\!\bra{m_{1},m_{2}}$.
[…]
$E_{2}(n_{1},n_{2})=E_{0}+4J\!\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\mathcal{N}_{\alpha}}\frac{1}{r^{\alpha}_{n_{1},n_{2}}}.$
Diagonal term.
[…]
$U(\mathbf{d})$ $=E_{0}+4J\!\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}+g^{2}\sum_{\beta\in(1,3)\text{–flip}}\frac{|\langle\mathbf{d}|\hat{T}|\beta\rangle|^{2}}{E_{2}(\mathbf{d})-E_{\beta}}$
$=E_{0}+4J\!\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}$
$\quad+\frac{g^{2}}{4}\left[\frac{2}{2J\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}}-\sum_{\begin{subarray}{c}\mathbf{d^{\prime}}\neq(0,0)\\ \mathbf{d^{\prime}}\neq\mathbf{d}\end{subarray}}\frac{1}{2J\left(1-\frac{1}{L^{2}}\right)-\frac{2J}{\mathcal{N}_{\alpha}}\left[\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}+\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}\right]}\right].$
Taking the large-$L$ limit gives the leading–order corrections. The first correction term can be written as
$\frac{g^{2}}{8J}\frac{2}{\left(1-\frac{1}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}\right)}$ $=\frac{g^{2}}{4J}\left[1-\frac{1}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}\right]^{-1}\approx\frac{g^{2}}{4J}\left[1+\frac{1}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}\right]=\frac{g^{2}}{4J}+\frac{g^{2}}{4J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}.$
The second correction term is
$\frac{g^{2}}{8J}\sum_{\begin{subarray}{c}\mathbf{d^{\prime}}\neq(0,0)\\ \mathbf{d^{\prime}}\neq\mathbf{d}\end{subarray}}\frac{1}{\left[1-\frac{1}{\mathcal{N}_{\alpha}}\left(\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}+\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}\right)\right]}$ $=\frac{g^{2}}{8J}\sum_{\begin{subarray}{c}\mathbf{d^{\prime}}\neq(0,0)\\ \mathbf{d^{\prime}}\neq\mathbf{d}\end{subarray}}\left[1+\sum^{\infty}_{i=1}\left(\frac{1}{\mathcal{N}_{\alpha}}\right)^{i}\left(\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}+\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}\right)^{i}\right]$
$=\frac{g^{2}(L^{2}-2)}{8J}+\frac{g^{2}}{8J}\sum^{\infty}_{i=1}\left[\left(\frac{1}{\mathcal{N}_{\alpha}}\right)^{i}\sum_{\begin{subarray}{c}\mathbf{d^{\prime}}\neq(0,0)\\ \mathbf{d^{\prime}}\neq\mathbf{d}\end{subarray}}\left(\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}+\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}\right)^{i}\right]$
$=\frac{g^{2}(L^{2}-2)}{8J}+\frac{g^{2}}{8J}\sum^{\infty}_{i=1}\left[\left(\frac{1}{\mathcal{N}_{\alpha}}\right)^{i}\sum_{m=0}^{i}\binom{i}{m}\sum_{\begin{subarray}{c}\mathbf{d^{\prime}}\neq(0,0)\\ \mathbf{d^{\prime}}\neq\mathbf{d}\end{subarray}}\frac{1}{|\mathbf{d^{\prime}}|^{m\alpha}|\mathbf{d}-\mathbf{d^{\prime}}|^{(i-m)\alpha}}\right].$
Separating the $\mathbf{d}$–dependent terms, we obtain
$\frac{g^{2}}{8J}\sum_{\begin{subarray}{c}\mathbf{d^{\prime}}\neq(0,0)\\ \mathbf{d^{\prime}}\neq\mathbf{d}\end{subarray}}\frac{1}{\left[1-\frac{1}{\mathcal{N}_{\alpha}}\left(\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}+\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}\right)\right]}$ $=\frac{g^{2}(L^{2}-2)}{8J}+\frac{g^{2}}{4J}\sum_{i=1}^{\infty}\left(\frac{1}{\mathcal{N}_{\alpha}}\right)^{i}\left(\sum_{\mathbf{d^{\prime}}\neq(0,0)}\frac{1}{|\mathbf{d^{\prime}}|^{i\alpha}}\right)$
$\quad-\frac{g^{2}}{4J}\sum_{i=1}^{\infty}\left(\frac{1}{\mathcal{N}_{\alpha}}\right)^{i}\frac{1}{|\mathbf{d}|^{i\alpha}}+\frac{g^{2}}{8J}\sum_{i=2}^{\infty}\left(\frac{1}{\mathcal{N}_{\alpha}}\right)^{i}\sum^{i-1}_{m=2}\mathcal{F}^{i}_{m}(\mathbf{d}),$
where
$\mathcal{F}^{i}_{m}(\mathbf{d})=\sum_{\begin{subarray}{c}\mathbf{d^{\prime}}\neq(0,0)\\ \mathbf{d^{\prime}}\neq\mathbf{d}\end{subarray}}\frac{1}{|\mathbf{d^{\prime}}|^{m\alpha}|\mathbf{d}-\mathbf{d^{\prime}}|^{(i-m)\alpha}}$
can be evaluated for large $|\mathbf{d}|$ by approximating the sum by a continuum integral,
$\mathcal{F}^{i}_{m}(R)=\int d^{2}r\,\frac{1}{|r|^{m\alpha}|r-R|^{(i-m)\alpha}}.$
Using the homogeneous function property, one finds
$\mathcal{F}^{i}_{m}(\lambda R)=\frac{1}{\lambda^{(i\alpha-2)}}\mathcal{F}^{i}_{m}(R)\quad\Rightarrow\quad\mathcal{F}^{i}_{m}(R)=\frac{\mathcal{F}^{i}_{m}(1)}{R^{(i\alpha-2)}}.$
Having all these ingredients, we can now write down the second correction term, ignoring higher–order contributions in $i$:
$\frac{g^{2}(L^{2}+2)}{8J}-\frac{g^{2}}{4J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}+\frac{g^{2}}{8J\mathcal{N}^{2}_{\alpha}}\frac{\mathcal{F}^{i=2}_{m=1}(1)}{|\mathbf{d}|^{2(\alpha-1)}}.$
Finally, combining everything gives the expression for $U(\mathbf{d})$:
$\tilde{U}(\mathbf{d})=E_{0}+4J-\frac{g^{2}L^{2}}{8J}-\frac{(2J-g^{2}/2J)}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}},$
and we isolate the $\mathbf{d}$–dependent potential as
$U(\mathbf{d})=-\frac{(2J-g^{2}/2J)}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}.$
Off–diagonal term.
The off–diagonal matrix elements between distinct separations are
[…]
$=\frac{g^{2}}{8}\left(\frac{1}{2J(1-1/L^{2})-\frac{2J}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}}+\frac{1}{2J(1-1/L^{2})-\frac{2J}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}}\right)$
$\quad-\frac{g^{2}}{8}\left(\frac{1}{2J(1-1/L^{2})-\frac{2J}{\mathcal{N}_{\alpha}}\left(\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}+\frac{1}{|\mathbf{d-d^{\prime}}|^{\alpha}}\right)}+\frac{1}{2J(1-1/L^{2})-\frac{2J}{\mathcal{N}_{\alpha}}\left(\frac{1}{|\mathbf{d}|^{\alpha}}+\frac{1}{|\mathbf{d-d^{\prime}}|^{\alpha}}\right)}\right).$
Taking the large-$L$ limit and expanding to first order in $\frac{1}{|\mathbf{d}|^{\alpha}}$, we get
$t_{\mathbf{d},\mathbf{d}^{\prime}}$ $\approx\frac{g^{2}}{16J}\Biggl(1+\frac{1}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}}+1+\frac{1}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}$
$\qquad\qquad-1-\frac{1}{\mathcal{N}_{\alpha}}\left(\frac{1}{|\mathbf{d}|^{\alpha}}+\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}\right)-1-\frac{1}{\mathcal{N}_{\alpha}}\left(\frac{1}{|\mathbf{d^{\prime}}|^{\alpha}}+\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}\right)\Biggr),$
which simplifies to
$t_{\mathbf{d},\mathbf{d}^{\prime}}\approx-\frac{g^{2}}{8J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}.$
Combining the diagonal and off–diagonal processes gives
[…]
with
$E_{2}$ $\approx E_{0}+4J-\frac{g^{2}L^{2}}{8J},\qquad U(\mathbf{d})\approx-\frac{(2J-g^{2}/2J)}{\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{\alpha}},\qquad t_{\mathbf{d},\mathbf{d}^{\prime}}\approx-\frac{g^{2}}{8J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}-\mathbf{d^{\prime}}|^{\alpha}}.$